OpenStudy (anonymous):

I will give a medal and fan anyone who can help explain this!!! A chocolate bar measures 40 mm wide, 80 mm long, and mm high. Will's Wrapping Company is making a wrapper to cover the chocolate bar. Use the correct net and determine how much paper will be needed to make the wrapper.

OpenStudy (anonymous):

https://lvs.brainhoney.com/Resource/45133303,0/Assets/84734_5508510b/05_07_20.jpg Will's Wrapping Company needs 3,860 mm2 of paper to wrap the chocolate bar. Will's Wrapping Company needs 7,720 mm2 of paper to wrap the chocolate bar. Will's Wrapping Company needs 8,160 mm2 of paper to wrap the chocolate bar. Will's Wrapping Company needs 15,440 mm2 of paper to wrap the chocolate bar. @Directrix @freckles @ILUVSOCCER @jdoe0001 @jigglypuff314 @mathmale @phi @pooja195 @satellite73 @YanaSidlinskiy

OpenStudy (anonymous):

The formula is A = 2(wl + hl + hw) Where w = width, h = height, and l = length. We know that w = 512mm l = 80mm h = 40mm So we have to plug in the values in the formula A = 2(512×80 + 40×80 + 40×512)

OpenStudy (anonymous):

We can convert 512 to 5.5 to make things easier A = 2(5.5×80 + 40×80 + 40×5.5)

OpenStudy (anonymous):

A = 2(5.5×80 + 40×80 + 40×5.5) We have to do what is int he parenthesis first, and in the parenthesis we have to do all the multiplication first starting from the left side of the equation A = 2(5.5×80 + 40×80 + 40×5.5)

OpenStudy (anonymous):

so 7720 right?

OpenStudy (anonymous):

Yes =) And that is your answer

OpenStudy (anonymous):

thanks so much =D

OpenStudy (anonymous):

youre welcome